OpenStudy (anonymous):

f(x) (1/5)to the x power x=0 simplify

OpenStudy (anonymous):

anyone?

myininaya (myininaya):

((1/5)^x)x=0?

myininaya (myininaya):

\[(\frac{1}{5})^x*x=0\]

myininaya (myininaya):

is this right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

x=0 and (1/5)^x=0 (1/5)^x is never 0 so the only solution is x=0